OpenStudy (anonymous):

Car 1 is travelling at 30m/s East when it collides with Car 2 who was travelling at 20m/s North. They stick together and continue to travel forwards, what is the final velocity of the clump?

OpenStudy (anonymous):

Im not sure how to do this because im not sure about the masses but i'll still try m1v1+m2v2= 2m3v3 The equation wont be is (-) beause they are not moving in opposite direction Considering the masses to be the same m1x30+m2x20=2m3v3 the mass gets cancelled as it is the same leaving 30+20= 2v3 50/2=v3 25=v3

OpenStudy (anonymous):

really sorry the masses are 3900 for car 1 and 1890 for car 2. is it possible now?

OpenStudy (anonymous):

There should be another value of veloctiy given

OpenStudy (anonymous):

they have the same final velocity, they are stuck together.

OpenStudy (rane):

this will help u how to find the final velocity

OpenStudy (rane):

after the collision

OpenStudy (anonymous):

Thanks :)

OpenStudy (rane):

yw:)